### v1

# The single edge notch fracture test for viscoelastic elastomers

Abstract
Making use of the Griffith criticality condition recently introduced by Shrimali and Lopez-Pamies (Extreme Mechanics Letters 58: 101944, 2023), this work presents a comprehensive analysis of the single edge notch fracture test for viscoelastic elastomers. The results — comprised of a combination of a parametric study and direct comparisons with experiments — reveal how the non-Gaussian elasticity, the nonlinear viscosity, and the intrinsic fracture energy of elastomers interact and govern when fracture nucleates from the pre-existing crack in these tests. The results also serve to quantify the limitations of existing analyses, wherein viscous effects and the actual geometries of the pre-existing cracks and the specimens are neglected.
Keywords:
Rubber; Elastomers; Adhesives; Cavitation; Fracture
1 Introduction
As a direct consequence of two elementary observations, Shrimali and Lopez-PamiesShrimali and Lopez-Pamies (2023a) have shown that the original formRivlin and Thomas (1953); Greensmith and Thomas (1955)
$-\dfrac{\partial\mathcal{W}}{\partial\mathrm{\Gamma}_{0}}=T_{c}$ (1)
of the Griffith criticality condition that describes the growth of cracks in viscoelastic elastomers [^1] By “viscoelastic elastomers” we mean elastomers that in addition to being capable to store energy by elastic deformation, they are capable to dissipate energy by viscous deformation. Dissipation mechanisms other than viscous deformation and fracture (e.g., Mullins effect, strain-induced crystallization) are considered to be negligible or absent altogether. subjected to quasi-static mechanical loads can be reduced to the fundamental form
$-\dfrac{\partial\mathcal{W}^{{\rm Eq}}}{\partial\mathrm{\Gamma}_{0}}=G_{c}.$ (2)
In expression (1), the left-hand side $-\partial\mathcal{W}/\partial\mathrm{\Gamma}_{0}$ denotes the change [^2] The derivative $-\partial\mathcal{W}/\partial\mathrm{\Gamma}_{0}$ is to be carried out under conditions of fixed deformation of those parts of the boundary that are not traction-free so that no work is done by the external forces. in total deformation — stored [^3] The stored part of the energy is comprised of an equilibrium part $\mathcal{W}^{{\rm Eq}}$ and a non-equilibrium part $\mathcal{W}^{{\rm NEq}}$, the latter representing the part of the stored energy that gets dissipated via viscous deformation eventually. and dissipated — energy
$\mathcal{W}=\underbrace{\mathcal{W}^{{\rm Eq}}+\mathcal{W}^{{\rm NEq}}}_{\text{stored}}+\underbrace{\mathcal{W}^{v}}_{\text{dissipated}}$
in the specimen at hand with respect to an added surface area ${\rm d}\mathrm{\Gamma}_{0}$ to the pre-existing crack $\mathrm{\Gamma}_{0}$ in its reference state, while the right-hand side $T_{c}$ stands for the critical tearing energy, a characteristic property of the elastomer that depends on the loading history. Traditionally, the critical tearing energy has been written in the form Gent and Schultz (1972); Gent (1996)
$T_{c}=G_{c}(1+f_{c}),$
where the material constant $G_{c}$ denotes the intrinsic fracture energy, or critical energy release rate, associated with the creation of new surface in the given elastomer, while $f_{c}$ stands for a non-negative material function of the loading history.
In expression (2), on the other hand, the left-hand side $-\partial\mathcal{W}^{{\rm Eq}}/$ $\partial\mathrm{\Gamma}_{0}$ denotes the change not in $\mathcal{W}$ but only in equilibrium elastic energy $\mathcal{W}^{{\rm Eq}}$ stored in the specimen with respect to an added surface area to the pre-existing crack $\mathrm{\Gamma}_{0}$, while the right-hand side involves not $T_{c}$ but only the intrinsic fracture energy $G_{c}$ of the elastomer.
From a fundamental point of view, the criticality condition (2) states that whether an elastomer deforms or creates new surface from a pre-existing crack is dictated by a competition solely between its stored equilibrium elastic energy and its intrinsic fracture energy. A corollary of this result is that the critical tearing energy $T_{c}$ in (1) — whose dependence via $f_{c}$ on the loading history had remained unknown for decades, preventing the use of the criticality condition (1) as a predicting theory — is given by the formula
$T_{c}=G_{c}(1+f_{c})=G_{c}-\dfrac{\partial\mathcal{W}^{{\rm NEq}}}{\partial\mathrm{\Gamma}_{0}}-\dfrac{\partial\mathcal{W}^{v}}{\partial\mathrm{\Gamma}_{0}},$ (3)
where the right-hand side is to be evaluated at the instance in time at which (2) is attained along the loading path of interest. This last relation clarifies how $T_{c}$ describes the total energy (per unit undeformed fracture area) expended in the tearing process, as it makes explicit the contributions by the creation of new surface ($G_{c}$) and by viscous effects ($-\partial\mathcal{W}^{{\rm NEq}}/\partial\mathrm{\Gamma}_{0}-\partial\mathcal{W}^{v}/\partial\mathrm{\Gamma}_{0}$).
From a practical point of view, the criticality condition (2) has the merit of being straightforward to employ for arbitrary specimen geometries and arbitrary loading conditions. This is because its evaluation requires having knowledge of the finite viscoelastic behavior of the elastomer — from which the stored equilibrium elastic energy $\mathcal{W}^{{\rm Eq}}$, as well as the stored non-equilibrium elastic energy $\mathcal{W}^{{\rm NEq}}$ and the dissipated viscous energy $\mathcal{W}^{v}$, can be computed — and of its intrinsic fracture energy. Both of these are material properties that can be measured once and for all.
Figure 1: Schematic of the single edge notch fracture test for a viscoelastic elastomer illustrating the geometry of the specimen, the applied time-dependent separation $l(t)$ between the grips, and the corresponding force $P(t)$ at the grips. It is standard practice to use the normalized quantities $\mathrm{\Lambda}=l(t)/L$ and $S=P(t)/(HB)$. We refer to them, respectively, as the global stretch and the global stress. Importantly, the initial size $A$ of the crack (the “notch”) can be equal to, larger, or smaller than the initial thickness $B$ of the specimen.
As a first application of the criticality condition (2), Shrimali and Lopez-PamiesShrimali and Lopez-Pamies (2023a) employed it to explain the “pure-shear” fracture test. In two ensuing companion papersShrimali and Lopez-Pamies (2023b, c), the same authors made use of (2) to explain two other types of prominent tests used to study fracture in viscoelastic elastomers: the delayed fracture test and the trousers fracture test.
The objective of this paper is to make use of the criticality condition (2) to explain in a detailed and quantitative manner yet another of the staple fracture tests utilized in the literature to probe the growth of cracks in elastomers, that depicted schematically in Fig. 1: the single edge notch fracture test.
1.1 A comment on previous works on the single edge notch fracture test
The first experimental results reported in the literature of single edge notch fracture tests applied to elastomers date back to the 1930s. They correspond to a series of tests carried out by BusseBusse (1934) with specimens made of natural rubber that featured a wide range of sizes for the pre-existing crack.
The early work by BusseBusse (1934) went unnoticed in the mechanics community until Rivlin and ThomasRivlin and Thomas (1953) presented their own single edge notch fracture tests alongside a first theoretical analysis in the 1950s. Their analysis amounts to the approximate evaluation of the criticality condition (1) making use not of local fields but of global quantities — in particular, the separation $l(t)$ between the grips and the corresponding force $P(t)$ at the grips — and of two simplifying constitutive and geometric assumptions:
  1. i.
    That the mechanical behavior of the elastomer that the specimen is made of is purely elastic; and
  2. ii.
    That the size $A$ of the pre-existing crack is much smaller than the length $L$ and the width $H$ of the specimen, but much larger than its thickness $B$ (see Fig. 1), so that boundary effects can be neglected and plane stress conditions apply.
Under these assumptions, which, as illustrated in Subsection 4.4 below, are rarely satisfied in practice, they concluded that
$-\dfrac{\partial\mathcal{W}}{\partial\mathrm{\Gamma}_{0}}=2AK(\mathrm{\Lambda})W_{{\rm un}}(\mathrm{\Lambda})=T_{c},$ (4)
where $W_{{\rm un}}(\mathrm{\Lambda})$ stands for the stored-energy function of the elastomer under uniaxial tension evaluated at the global stretch $\mathrm{\Lambda}=l/L$, while $K(\mathrm{\Lambda})$ is a numerical factor that varies with $\mathrm{\Lambda}$.
In 1963, a decade after Rivlin and ThomasRivlin and Thomas (1953) introduced the result (4), GreensmithGreensmith (1963) carried out single edge notch fracture tests on four different types of natural rubbers aimed at determining the numerical factor $K(\mathrm{\Lambda})$ in (4) directly from experimental measurements. The fitting of (4) to his data indicated that, irrespective of the type of rubber that he tested, $K(\mathrm{\Lambda})$ decreases nonlinearly from a value of about $3$ at small stretches when $\mathrm{\Lambda}\approx 1$ to a value somewhat below $2$ at $\mathrm{\Lambda}\approx 3$. Later, in 1972, guided by the experiments of GreensmithGreensmith (1963) and by his own 2D plane stress finite element (FE) simulations of the test, LindleyLindley (1972) proposed the approximation
$K(\mathrm{\Lambda})=\dfrac{2.95-0.08(\mathrm{\Lambda}-1)}{\sqrt{\mathrm{\Lambda}}}\approx\dfrac{3}{\sqrt{\mathrm{\Lambda}}}.$ (5)
By now, experiments beyond those of GreensmithGreensmith (1963) are abundant [^4] The single edge notch fracture test has also been of recent popular use to study fracture in elastomers containing a liquid solventKwon et al. (2011); Guo et al. (2016); Lin et al. (2022). in the literatureHamed and Park (1999); Cristiano et al. (2011); Fan et al. (2017); Roucou et al. (2018); Slootman et al. (2020); Menn (2022); Wang and Fan (2023); Xue et al. (2024). However, most of these experiments have been carried out at a fixed constant stretch rate — wherein a global stretch of the form $\mathrm{\Lambda}=1+\dot{\mathrm{\Lambda}}_{0}t$, with fixed $\dot{\mathrm{\Lambda}}_{0}$, is prescribed — which conceals the viscous dissipation that inevitably takes place in these tests. Furthermore, many of these experiments have been carried out with specimens for which the size of the pre-existing crack is comparable to both the specimen width and the specimen thickness, that is, $A\sim H$ and $A\sim B$, so that boundary effects and the 3D character of the mechanical fields are not expected to be negligible.
Beyond the simulations of LindleyLindley (1972), a plurality of FE simulations are also now available in the literatureYeoh (2002); Chen et al. (2017); Kang et al. (2020); Liu et al. (2020). Invariably, all of these simulations assume that the mechanical behavior of elastomers is purely elastic. What is more, they are restricted to 2D plane stress simulations.
In this context, this work aims at reporting an analysis that fully accounts for the viscoelasticity of elastomers and the 3D geometry of the pre-existing crack and the specimen in single edge notch fracture tests, and hence one that should allow to explain and appropriately interpret the results from such tests.
2 Formulation of the initial-boundary-value problem for the single edge notch fracture test
2.1 Initial configuration and kinematics
Consider the rectangular specimens depicted in Fig. 1 of length $L$ and width $H$ in the ${\bf e}_{3}$ and ${\bf e}_{2}$ directions and constant thickness $B<H$ in the ${\bf e}_{1}$ direction. The specimens contain a pre-existing edge crack of length $A<H$ in the ${\bf e}_{2}$ direction. Here, $\{{\bf e}_{i}\}$ $(i=1,2,3)$ stands for the laboratory frame of reference. We place its origin at the specimens’ midplane along the edge containing the crack so that, in their initial configuration at time $t=0$, the specimens occupy the domain
$\overline{\mathrm{\Omega}}_{0}=\{{\bf X}:{\bf X}\in\mathcal{P}_{0}\setminus\mathrm{\Gamma}_{0}\},$
where
[…]
and
[…]
At a later time $t\in(0,T]$, in response to the boundary conditions described below, the position vector ${\bf X}$ of a material point occupies a new position ${\bf x}$ specified by an invertible mapping ${\bf y}$ from $\mathrm{\Omega}_{0}$ to the current configuration $\mathrm{\Omega}(t)\subset\mathbb{R}^{3}$. We write
${\bf x}={\bf y}({\bf X},t)$
and the associated deformation gradient and velocity fields at ${\bf X}\in\mathrm{\Omega}_{0}$ and $t\in(0,T]$ as
${\bf F}({\bf X},t)=\nabla{\bf y}({\bf X},t)=\frac{\partial{\bf y}}{\partial{\bf X}}({\bf X},t)$
and
$\textbf{V}({\bf X},t)=\dot{{\bf y}}({\bf X},t)=\frac{\partial{\bf y}}{\partial t}({\bf X},t).$
Throughout, we shall use the “dot” notation to denote the material time derivative (i.e., with ${\bf X}$ held fixed) of field quantities.
2.2 Constitutive behavior of the elastomer
The specimens are taken to be made of an isotropic incompressible elastomer. Making use of the two-potential formalism Kumar and Lopez-Pamies (2016), we describe its constitutive behavior by two thermodynamic potentials, a free energy density, or stored-energy function, of the form
$\psi({\bf F},{\bf C}^{v})=\left\{\begin{array}[]{ll}\psi^{{\rm Eq}}(I_{1})+\psi^{{\rm NEq}}(I_{1}^{e})&\,{\rm if}\,\,J=1\\ \\ +\infty&\,{\rm else}\end{array}\right.$ (6)
that describes how the elastomer stores energy through elastic deformation and a dissipation potential of the form
$\phi({\bf F},{\bf C}^{v},\dot{{\bf C}}^{v})=\left\{\begin{array}[]{ll}\dfrac{1}{2}\dot{{\bf C}}^{v}\cdot\mbox{\boldmath$\mathcal{A}$}({\bf F},{\bf C}^{v})\dot{{\bf C}}^{v}&\,{\rm if}\,\,J^{v}=1\\ \\ +\infty&\,{\rm else}\end{array}\right.$ (7)
with $\mathcal{A}_{ijkl}({\bf F},{\bf C}^{v})=\dfrac{\eta({\bf C},{\bf C}^{v})}{2}{C^{v}}^{-1}_{ik}{C^{v}}^{-1}_{jl}$ that describes how the elastomer dissipates energy through viscous deformation. In these expressions, the symmetric second-order tensor ${\bf C}^{v}$ is an internal variable of state that stands for a measure of the “viscous part” of the deformation gradient ${\bf F}$,
$I_{1}={\rm tr}\,{\bf C},\quad J=\sqrt{\det{\bf C}},\quad I_{1}^{e}={\rm tr}({\bf C}{{\bf C}^{v}}^{-1}),\quad J^{v}=\sqrt{\det{\bf C}^{v}},$
where ${\bf C}={\bf F}^{T}{\bf F}$ denotes the right Cauchy-Green deformation tensor, and $\psi^{{\rm Eq}}$, $\psi^{{\rm NEq}}$, $\eta$ are any suitably well-behaved non-negative material functions of their arguments.
Granted the two thermodynamic potentials (6) and (7), it follows that the first Piola-Kirchhoff stress tensor ${\bf S}$ at any material point ${\bf X}\in\mathrm{\Omega}_{0}$ and time $t\in[0,T]$ is given by the relation Kumar and Lopez-Pamies (2016)
${\bf S}({\bf X},t)=\frac{\partial\psi}{\partial{\bf F}}({\bf F},{\bf C}^{v}),$
where ${\bf C}^{v}$ is implicitly defined by the evolution equation
$\left\{\begin{array}[]{l}\dfrac{\partial\psi}{\partial{\bf C}^{v}}({\bf F},{\bf C}^{v})+\dfrac{\partial\phi}{\partial\dot{{\bf C}}^{v}}({\bf F},{\bf C}^{v},\dot{{\bf C}}^{v})={\bf 0}\\ \\ {\bf C}^{v}({\bf X},0)={\bf I}\end{array}\right..$
Making use of the specific isotropic incompressible forms (6) and (7), this relation can be rewritten more explicitly as
${\bf S}({\bf X},t)=2\psi^{{\rm Eq}}_{I_{1}}{\bf F}+2\psi^{{\rm NEq}}_{I^{e}_{1}}{\bf F}{{\bf C}^{v}}^{-1}-p{\bf F}^{-T},$ (8)
where $p$ stands for the arbitrary hydrostatic pressure associated with the incompressibility constraint $J=1$ of the elastomer and ${\bf C}^{v}$ is defined implicitly as the solution of the evolution equation
$\left\{\begin{array}[]{l}\dot{{\bf C}}^{v}({\bf X},t)=\dfrac{2\psi^{{\rm NEq}}_{I^{e}_{1}}}{\eta({\bf C},{\bf C}^{v})}\left({\bf C}-\dfrac{1}{3}\left({\bf C}\cdot{{\bf C}^{v}}^{-1}\right){\bf C}^{v}\right)\\ \\ {\bf C}^{v}({\bf X},0)={\bf I}\end{array}\right..$ (9)
In these last expressions, we have made use of the notation $f_{x}={\rm d}f(x)/{\rm d}x$ for the derivative of scalar-valued functions of one variable.
For a detailed account of the constitutive relation (8)-(9) and its physical interpretation, the interested reader is referred to Kumar and Lopez-PamiesKumar and Lopez-Pamies (2016). Here, it suffices to make the following remark.
Remark 1
The constitutive relation (8)-(9) is nothing more than a generalization of the classical Zener or standard solid model Zener (1948) to the setting of finite deformations. To see this, note that the rheological representation of the constitutive relation (8)-(9) is given by the Zener rheological representation shown in Fig. 2. Note as well that in the limit of small deformations, as ${\bf F}\rightarrow{\bf I}$, the constitutive relation (8)-(9) reduces to the classical linear viscoelastic Zener model
${\bf S}({\bf X},t)=2\,\mu^{{\rm Eq}}\,\boldsymbol{\varepsilon}+2\,\mu^{{\rm NEq}}(\boldsymbol{\varepsilon}-\boldsymbol{\varepsilon}^{v})-p{\bf I},$
where $\boldsymbol{\varepsilon}=1/2({\bf F}+{\bf F}^{T}-2{\bf I})$ stands for the infinitesimal strain tensor and $\boldsymbol{\varepsilon}^{v}=1/2({\bf C}^{v}-{\bf I})$ is solution of the evolution equation
$\left\{\begin{array}[]{l}\dot{\boldsymbol{\varepsilon}}^{v}({\bf X},t)=\dfrac{\mu^{{\rm NEq}}}{\eta_{0}}\left(\boldsymbol{\varepsilon}-\boldsymbol{\varepsilon}^{v}-\dfrac{1}{3}{\rm tr}(\boldsymbol{\varepsilon}-\boldsymbol{\varepsilon}^{v}){\bf I}\right)\\ \\ \boldsymbol{\varepsilon}^{v}({\bf X},0)=\textbf{0}\end{array}\right..$
In these expressions, $\mu^{{\rm Eq}}=2\psi_{I_{1}}^{{\rm Eq}}(3)$ and $\mu^{{\rm NEq}}=2\psi_{I^{e}_{1}}^{{\rm NEq}}(3)$ stand for the initial shear moduli associated with the equilibrium and non-equilibrium stored-energy functions, respectively, while $\eta_{0}=\eta({\bf I},{\bf I})$ stands for the initial viscosity.
Figure 2: Rheological representation of the two-potential viscoelastic model (8)-(9).
In the parametric study that is presented below, we make use of the following specific forms for the equilibrium and non-equilibrium stored-energy functions in (6) and viscosity function in (7):
$\left\{\begin{array}[]{l}\psi^{{\rm Eq}}(I_{1})=\displaystyle\sum\limits_{r=1}^{2}\dfrac{3^{1-\alpha_{r}}}{2\alpha_{r}}\mu_{r}\left(\,I_{1}^{\alpha_{r}}-3^{\alpha_{r}}\right)\vspace{0.2cm}\\ \psi^{{\rm NEq}}(I^{e}_{1})=\displaystyle\sum\limits_{r=1}^{2}\dfrac{3^{1-\beta_{r}}}{2\beta_{r}}\nu_{r}\left(\,{I^{e}_{1}}^{\beta_{r}}-3^{\beta_{r}}\right)\vspace{0.2cm}\\ \eta({\bf C},{\bf C}^{v})=\widetilde{\eta}(I_{1}^{v},\mathcal{J}_{2}^{{\rm NEq}})\end{array}\right.$ (10)
with $I_{1}^{v}={\rm tr}\,{\bf C}^{v}$ and $\mathcal{J}_{2}^{{\rm NEq}}=\left(\dfrac{{I_{1}^{e}}^{2}}{3}-I_{2}^{e}\right)\left(\displaystyle\sum\limits_{r=1}^{2}3^{1-\beta_{r}}\nu_{r}{I_{1}^{e}}^{\beta_{r}-1}\right)^{2}$, where $I_{2}^{e}=1/2({I_{1}^{e}}^{2}-{\rm tr}\,({\bf C}{{\bf C}^{v}}^{-1}{\bf C}{{\bf C}^{v}}^{-1}))$, which result in the constitutive relation
${\bf S}({\bf X},t)=$ $\displaystyle\sum\limits_{r=1}^{2}3^{1-\alpha_{r}}\mu_{r}I_{1}^{\alpha_{r}-1}{\bf F}+\displaystyle\sum\limits_{r=1}^{2}3^{1-\beta_{r}}\nu_{r}{I^{e}_{1}}^{\beta_{r}-1}{\bf F}{{\bf C}^{v}}^{-1}-p{\bf F}^{-T}$ (11)
with evolution equation
$\left\{\begin{array}[]{l}\dot{{\bf C}}^{v}({\bf X},t)=\dfrac{\sum\limits_{r=1}^{N}3^{1-\beta_{r}}\nu_{r}{I^{e}_{1}}^{\beta_{r}-1}}{\eta({\bf C},{\bf C}^{v})}\left({\bf C}-\dfrac{1}{3}\left({\bf C}\cdot{{\bf C}^{v}}^{-1}\right){\bf C}^{v}\right)\\ \\ {\bf C}^{v}({\bf X},0)={\bf I}\end{array}\right..$ (12)
The constitutive prescription (11)-(12) includes several fundamental constitutive relations as special cases. In particular, it includes the cases of:
  • $\bullet$
    A Neo-Hookean solid ($\mu_{2}=\nu_{1}=\nu_{2}=0$, $\alpha_{1}=1$, $\eta=0$);
  • $\bullet$
    A Newtonian fluid ($\mu_{1}=\mu_{2}=\nu_{2}=0$, $\nu_{1}=+\infty$, $\eta=\eta_{0}$);
  • $\bullet$
    As well as of a viscoelastic elastomer with Gaussian elasticity and constant viscosity ($\mu_{2}=\nu_{2}=0$, $\alpha_{1}=\beta_{1}=1$, $\eta=\eta_{0}$).
The prescription (11)-(12) has the additional merit of being able to describe the viscoelastic behavior of a wide range of elastomers, which typically exhibit non-Gaussian elasticityTreloar (2005); Lopez-Pamies (2010), as well as a strongly nonlinear viscosity that depends both on the deformation and on the deformation rateDoi and Edwards (1998); Bergström and Boyce (1998); Amin et al. (2006); Kumar and Lopez-Pamies (2016); Jianyou et al. (2018); Ghosh et al. (2021); Chockalingam et al. (2021); Chen and Ravi-Chandar (2022); Ricker et al. (2023). In particular, the viscosity of elastomers typically increases with increasing deformation and decreases with increasing deformation rate, the latter behavior being commonly referred to as shear-thinning. The viscosity function $\eta({\bf C},{\bf C}^{v})=\widetilde{\eta}(I_{1}^{v},\mathcal{J}_{2}^{{\rm NEq}})$ in (10) can be suitably prescribed to describe both of these behaviors via its dependence on deformation through $I_{1}^{v}$ and on deformation rate through $\mathcal{J}_{2}^{{\rm NEq}}$. In the parametric study that is presented next, we will use the specific constitutive prescriptionKumar and Lopez-Pamies (2016); Ghosh et al. (2021)
$\widetilde{\eta}(I_{1}^{v},\mathcal{J}_{2}^{{\rm NEq}})=\widetilde{\eta}_{\infty}(I^{v}_{1})+\dfrac{\widetilde{\eta}_{0}(I^{v}_{1})-\widetilde{\eta}_{\infty}(I^{v}_{1})}{1+\left(K_{2}\mathcal{J}_{2}^{{\rm NEq}}\right)^{\gamma_{2}}}$ (13)
with $\widetilde{\eta}_{0}(I^{v}_{1})=\eta_{0}+(N_{0}-\eta_{0})[\tanh\left(I_{1}^{v\gamma_{1}}-3^{\gamma_{1}}-K_{1}\right)+\tanh K_{1}]/(1+\tanh K_{1})$ and $\widetilde{\eta}_{\infty}(I^{v}_{1})=\eta_{\infty}$. In total, the constitutive prescription (11)-(12) with (13) contains $15$ material constants. Four of them, $\mu_{r}$ and $\alpha_{r}$ ($r=1,2$), serve to characterize the non-Gaussian elasticity of the elastomer at states of thermodynamic equilibrium. Another four, $\nu_{r}$ and $\beta_{r}$ ($r=1,2$), characterize the non-Gaussian elasticity at non-equilibrium states. Finally, the last seven constants, $\eta_{0}$, $N_{0}$, $\eta_{\infty}$, $K_{1}$, $\gamma_{1}$, $K_{2}$, $\gamma_{2}$, serve to characterize the nonlinear deformation-dependent shear-thinning viscosity.
2.3 Initial and boundary conditions
In the initial configuration $\mathrm{\Omega}_{0}$, the elastomer is presumed to be undeformed and stress free. This implies the initial conditions
$\left\{\begin{array}[]{l}{\bf y}({\bf X},0)={\bf X\vspace{0.2cm}}\\ p({\bf X},0)=\mu^{{\rm Eq}}+\mu^{{\rm NEq}}\vspace{0.2cm}\\ {\bf C}^{v}({\bf X},0)={\bf I}\end{array}\right.,\quad{\bf X}\in\overline{\mathrm{\Omega}}_{0},$
for the deformation field ${\bf y}({\bf X},t)$, the pressure field $p({\bf X},t)$, and the internal variable ${\bf C}^{v}({\bf X},t)$.
The bottom boundary
[…]
and the top boundary
[…]
of the specimens are held firmly by stiff grips that are pulled apart in the ±$\pm{\bf e}_{3}$ directions resulting in a separation $l(t)$ between the grips. Consistent with the manner in which the vast majority of experiments are carried out, we restrict attention to the case when the grips are pulled apart at a constant rate $\dot{l}_{0}$, that is, when the separation between the grips is given by $l(t)=L+\dot{l}_{0}t$. This results in tests carried out at the constant global stretch rate $\dot{\mathrm{\Lambda}}_{0}=\dot{l}_{0}/L$. Therefore, making use of the notation $\partial\mathrm{\Omega}^{\mathcal{N}}_{0}=\partial\mathrm{\Omega}_{0}\setminus\left(\partial\mathrm{\Omega}^{\mathcal{B}}_{0}\cup\partial\mathrm{\Omega}^{\mathcal{T}}_{0}\right)$ and $\textbf{s}({\bf X},t)={\bf S}{\bf N}$, where ${\bf N}$ stands for the outward unit normal to the boundary $\partial\mathrm{\Omega}_{0}$, we have the boundary conditions
[…]
with
$\left\{\begin{array}[]{l}\overline{y}^{\mathcal{B}}_{1}({\bf X},t)=X_{1}\vspace{0.15cm}\\ \overline{y}^{\mathcal{B}}_{2}({\bf X},t)=X_{2}\vspace{0.15cm}\\ \overline{y}^{\mathcal{B}}_{3}({\bf X},t)=X_{3}-\dfrac{\dot{l}_{0}}{2}t\end{array}\right.\quad{\rm and}\quad\left\{\begin{array}[]{l}\overline{y}^{\mathcal{T}}_{1}({\bf X},t)=X_{1}\vspace{0.15cm}\\ \overline{y}^{\mathcal{T}}_{2}({\bf X},t)=X_{2}\vspace{0.15cm}\\ \overline{y}^{\mathcal{T}}_{3}({\bf X},t)=X_{3}+\dfrac{\dot{l}_{0}}{2}t\end{array}\right..$
2.4 Governing equations
Neglecting inertia and body forces, the equations of balance of linear and angular momenta read as ${\rm Div}\,{\bf S}={\bf 0}$ and ${\bf S}{\bf F}^{T}={\bf F}{\bf S}^{T}$ for $({\bf X},t)\in\mathrm{\Omega}_{0}\times[0,T]$. The latter is automatically satisfiedKumar and Lopez-Pamies (2016); Sadik and Yavari (2024) by virtue of the objectivity of the thermodynamic potentials (6) and (7) and so, upon combining all the above ingredients together and introducing the notation
${\bf S}({\bf X},t)=2\psi^{{\rm Eq}}_{I_{1}}\nabla{\bf y}+2\psi^{{\rm NEq}}_{I^{e}_{1}}\nabla{\bf y}{{\bf C}^{v}}^{-1}-p\nabla{\bf y}^{-T}$
and
$\tau({\bf X},t)=\dfrac{\widetilde{\eta}(I_{1}^{v},\mathcal{J}_{2}^{{\rm NEq}})}{2\psi^{{\rm NEq}}_{I^{e}_{1}}}$
for the stress and the relaxation time function, the governing equations that describe the mechanical response of the specimens amount to the initial-boundary-value problem
$\left\{\begin{array}[]{ll}{\rm Div}\,{\bf S}={\bf 0},&({\bf X},t)\in\mathrm{\Omega}_{0}\times[0,T]\\ \\ \det\nabla{\bf y}=1,&({\bf X},t)\in\mathrm{\Omega}_{0}\times[0,T]\\ \\ {\bf y}({\bf X},t)=\overline{{\bf y}}^{\mathcal{B}}({\bf X},t),&({\bf X},t)\in\partial\mathrm{\Omega}^{\mathcal{B}}_{0}\times[0,T]\\ \\ {\bf y}({\bf X},t)=\overline{{\bf y}}^{\mathcal{T}}({\bf X},t),&({\bf X},t)\in\partial\mathrm{\Omega}^{\mathcal{T}}_{0}\times[0,T]\\ \\ {\bf S}{\bf N}=\textbf{0},&({\bf X},t)\in\partial\mathrm{\Omega}^{\mathcal{N}}_{0}\times[0,T]\\ \\ {\bf y}({\bf X},0)={\bf X},&{\bf X}\in\mathrm{\Omega}_{0}\\ \\ p({\bf X},0)=\mu^{{\rm Eq}}+\mu^{{\rm NEq}},&{\bf X}\in\mathrm{\Omega}_{0}\end{array}\right.$ (14)
coupled with
$\left\{\begin{array}[]{ll}\dot{{\bf C}}^{v}({\bf X},t)=\dfrac{1}{\tau}\left(\nabla{\bf y}^{T}\nabla{\bf y}-\dfrac{1}{3}\left(\nabla{\bf y}^{T}\nabla{\bf y}\cdot{{\bf C}^{v}}^{-1}\right){\bf C}^{v}\right),&\\ \\ &({\bf X},t)\in\mathrm{\Omega}_{0}\times[0,T]\\ \\ {\bf C}^{v}({\bf X},0)={\bf I},&{\bf X}\in\mathrm{\Omega}_{0}\end{array}\right.$ (15)
for the deformation field ${\bf y}({\bf X},t)$, the pressure field $p({\bf X},t)$, and the internal variable ${\bf C}^{v}({\bf X},t)$.
All the results that we present below are generated by solving numerically the initial-boundary-value problem (14)-(15) via a variant of the scheme introduced by Ghosh et al.Ghosh et al. (2021), which is based on a Crouzeix-Raviart FE discretization of space and a high-order explicit Runge-Kutta finite difference discretization of time, as well as by a related scheme that has been recently implemented as an Abaqus UMAT subroutineLefèvre et al. ; we also generated corresponding 2D plane-stress results, which turned out to be in agreement with the fully 3D results. FEniCSx and Abaqus implementations of these numerical schemes are available on GitHub. [^1] http://pamies.cee.illinois.edu/Subroutines.html.
3 The Griffith criticality condition (2) for the single edge notch fracture test
Having established the governing equations (14)-(15) that describe the pointwise mechanical fields in the single edge notch fracture test for viscoelastic elastomers, the evaluation of the Griffith criticality condition (2) to determine when the pre-existing crack grows in such a test is straightforward. The pertinent calculations amount to the following three successive steps:
  1. i.
    Solving the equations (14)-(15) for the deformation field ${\bf y}({\bf X},t)$, the pressure field $p({\bf X},t)$, and the internal variable ${\bf C}^{v}({\bf X},t)$ for the material constants ($\mu_{r},\alpha_{r}$, $\nu_{r},\beta_{r}$ $\eta_{0}$, $N_{0}$, $\eta_{\infty}$, $K_{1}$, $\gamma_{1}$, $K_{2}$, $\gamma_{2}$) of choice, the specimen geometry ($L,H,B$) of choice, and sets of global stretch rates and sizes of the pre-existing crack that contain the specific values $\dot{\mathrm{\Lambda}}_{0}$ and $A$ of interest;
  2. ii.
    Computing the resulting deformation gradient ${\bf F}({\bf X},t)=\nabla{\bf y}$ and associated first invariant $I_{1}({\bf X},t)={\rm tr}\,{\bf F}^{T}{\bf F}$, evaluating the equilibrium energy
    $\mathcal{W}^{{\rm Eq}}(\dot{\mathrm{\Lambda}}_{0},A;\mathrm{\Lambda})=\displaystyle\int_{\mathrm{\Omega}_{0}}\psi^{{\rm Eq}}(I_{1})\,{\rm d}{\bf X},$ (16)
    and viewing the result as a function of the global stretch rate $\dot{\mathrm{\Lambda}}_{0}$ and the size $A$ of the pre-existing crack, parameterized by the global stretch $\mathrm{\Lambda}$; and
  3. iii.
    Computing the derivative
    $-\dfrac{\partial\mathcal{W}^{{\rm Eq}}}{\partial{\mathrm{\Gamma}_{0}}}=-\dfrac{1}{B}\dfrac{\partial\mathcal{W}^{{\rm Eq}}}{\partial{A}}(\dot{\mathrm{\Lambda}}_{0},A;\mathrm{\Lambda})$ (17)
    of the equilibrium energy so as to finally determine from the Griffith criticality condition
    $-\dfrac{1}{B}\dfrac{\partial\mathcal{W}^{{\rm Eq}}}{\partial{A}}(\dot{\mathrm{\Lambda}}_{0},A;\mathrm{\Lambda}_{c})=G_{c}$ (18)
    the critical value $\mathrm{\Lambda}_{c}$ of the global stretch $\mathrm{\Lambda}$ at which the crack grows for the global stretch rate $\dot{\mathrm{\Lambda}}_{0}$, initial crack size $A$, and critical energy release rate $G_{c}$ of interest.
Remark 2
Here, it is important to emphasize that the equilibrium energy (16) — and hence its derivative (17) — depends on the entire viscoelastic behavior of the elastomer through the solution of the governing equations (14)-(15) for the deformation field ${\bf y}({\bf X},t)$. Accordingly, the Griffith criticality condition (18) is expected to depend (possibly strongly) on the loading rate $\dot{\mathrm{\Lambda}}_{0}$.
Note that once the global stretch $\mathrm{\Lambda}_{c}$ at which the crack grows for a given test has been determined from the Griffith criticality condition (18), it is a simple matter to determine the associated tearing energy (3). The result can be written as
$T_{c}=G_{c}-\dfrac{1}{B}\dfrac{\partial\mathcal{W}^{{\rm NEq}}}{\partial{A}}(\dot{\mathrm{\Lambda}}_{0},A;\mathrm{\Lambda}_{c})-\dfrac{1}{B}\dfrac{\partial\mathcal{W}^{v}}{\partial{A}}(\dot{\mathrm{\Lambda}}_{0},A;\mathrm{\Lambda}_{c})$
with
$\mathcal{W}^{{\rm NEq}}(\dot{\mathrm{\Lambda}}_{0},A;\mathrm{\Lambda})=\displaystyle\int_{\mathrm{\Omega}_{0}}\psi^{{\rm NEq}}(I^{e}_{1})\,{\rm d}{\bf X}$ (19)
and
$\mathcal{W}^{v}(\dot{\mathrm{\Lambda}}_{0},A;\mathrm{\Lambda})=\displaystyle\int_{\mathrm{\Omega}_{0}}\displaystyle\int_{0}^{t}2\phi({\bf F}({\bf X},\tau),{\bf C}^{v}({\bf X},\tau),\dot{{\bf C}}^{v}({\bf X},\tau))\,{\rm d}\tau\,{\rm d}{\bf X},$ (20)
where the integrals in (19) and (20) are to be computed — much like that in (16) — in terms of the numerical solution of equations (14)-(15). Alternatively,
$T_{c}=-\dfrac{1}{B}\dfrac{\partial\mathcal{W}}{\partial{A}}(\dot{\mathrm{\Lambda}}_{0},A;\mathrm{\Lambda}_{c}),$ (21)
where
$\mathcal{W}(\dot{\mathrm{\Lambda}}_{0},A;\mathrm{\Lambda})=\displaystyle\int_{L}^{\mathrm{\Lambda}L}P\,{\rm d}l$
stands for the total deformation (stored and dissipated) energy, written here in terms of the area under the force-deformation ($P$ vs. $l$) response of the specimen.
Figure 3: Representative FE solution for the equilibrium stored-energy function $\psi^{{\rm Eq}}(I_{1})$ (shown in MPa) over the initial $\mathrm{\Omega}_{0}$ and current $\mathrm{\Omega}(t)$ configurations of a specimen, with pre-existing crack of length $A=2$ mm, at a global stretch $\mathrm{\Lambda}=1.5$, that has been stretched at the global stretch rate $\dot{\mathrm{\Lambda}}_{0}=10^{0}$ s-1.
Figure 4: Representative examples of the equilibrium energy (16) and its derivative (17) plotted as functions of the global stretch rate $\dot{\mathrm{\Lambda}}_{0}$ and the initial crack size $A$ at a fixed global stretch $\mathrm{\Lambda}$.
Figure 5: Representative example of the derivative (17) of the equilibrium elastic energy plotted as a function of the global stretch $\mathrm{\Lambda}$ and the global stretch rate $\dot{\mathrm{\Lambda}}_{0}$ for a fixed initial crack size $A$.
By way of an example, Fig. 3 provides a representative result for the equilibrium stored-energy function $\psi^{{\rm Eq}}(I_{1})$ over the initial $\mathrm{\Omega}_{0}$ and the current $\mathrm{\Omega}(t)$ configurations of a specimen, with pre-existing crack of length $A=2$ mm, at a global stretch $\mathrm{\Lambda}=1.5$, that has been stretched at the global stretch rate $\dot{\mathrm{\Lambda}}_{0}=10^{0}$ s-1; see also Fig. S1 in the Supplementary Material. Figure 4 presents corresponding results for the equilibrium energy (16) and its derivative (17) plotted as functions of the global stretch rate $\dot{\mathrm{\Lambda}}_{0}$ and the initial crack size $A$ at a fixed global stretch $\mathrm{\Lambda}$. Finally, Fig. 5 presents the derivative (17) of the equilibrium energy as a function of the global stretch $\mathrm{\Lambda}$ and the global stretch rate $\dot{\mathrm{\Lambda}}_{0}$ for a fixed initial crack size $A$. The results in Figs. 3 through 5 pertain to one of the single edge notch fracture tests considered in Section 4, that of an elastomer with material constants $\mu_{1}=0.09$ MPa, $\alpha_{1}=0.5$, $\mu_{2}=0.01$ MPa, $\alpha_{2}=2.5$, $\nu_{1}=1.8$ MPa, $\beta_{1}=0.5$, $\nu_{2}=0.2$ MPa, $\beta_{2}=7.5$, $\eta_{0}=N_{0}=\eta_{\infty}=2$ MPa s, and specimens of length $L=15$ mm, width $H=4$ mm, and thickness $B=0.5$ mm.
To illustrate the evaluation of the Griffith criticality condition (18), Fig. 5 includes a plane of constant value — in this case, $G_{c}=150$ N/m — whose intersection with the surface of the derivative $-\partial\mathcal{W}^{{\rm Eq}}/\partial{\mathrm{\Gamma}_{0}}$ allows to readily identify the critical global stretches $\mathrm{\Lambda}_{c}$ at which the crack grows for a given critical energy release rate $G_{c}$ as a function of the global stretch rate $\dot{\mathrm{\Lambda}}_{0}$.
4 A comprehensive parametric study of the effects of the elasticity and viscosity of the elastomer and of the 3D geometry of the cracks and specimens
At this stage, we are in a position to deploy the Griffith criticality condition (2) to analyze the single edge notch fracture test fully accounting for the viscoelasticity of elastomers and the 3D geometry of the pre-existing cracks and the specimens.
For definiteness, in the parametric study that follows, we consider specimens of length $L=15$ mm, width $H=4$ mm, and thickness $B=0.5$ mm that contain a pre-existing crack of three different lengths
$A=1.9,2,2.1\;\rm{mm}.$
These specific values for $L,H,B,A$ are chosen here because they are representative of those used in experimentsHamed and Park (1999); Cristiano et al. (2011); Fan et al. (2017); Roucou et al. (2018); Slootman et al. (2020); Menn (2022); Wang and Fan (2023); Xue et al. (2024).
As also representative of the behavior of typical elastomersLopez-Pamies (2010); Breedlove et al. (2024), we make use of the material constants
$\mu_{1}=0.09\,{\rm MPa},\quad\alpha_{1}=0.5,\quad\mu_{2}=0.01\,{\rm MPa},\quad\alpha_{2}=2.5$ (22)
in the equilibrium stored-energy function (10)1 that describe the non-Gaussian elasticity of the elastomer at states of thermodynamic equilibrium. Note that $\mu^{{\rm Eq}}=\mu_{1}+\mu_{2}=0.1$ MPa.
With the objective of thoroughly probing the effects of the additional non-Gaussian elasticity that arises at non-equilibrium states, we consider the three different sets of material constants
$\left\{\begin{array}[]{rl}\texttt{Weaker Growth:}&\left\{\begin{array}[]{l}\nu_{1}=0.9\nu,\;\beta_{1}=0.5\\ \nu_{2}=0.1\nu,\;\beta_{2}=1\end{array}\right.\vspace{0.18cm}\\ \texttt{Equal Growth:}&\left\{\begin{array}[]{l}\nu_{1}=0.9\nu,\;\beta_{1}=0.5\\ \nu_{2}=0.1\nu,\;\beta_{2}=2.5\end{array}\right.\vspace{0.18cm}\\ \texttt{Stronger Growth:}&\left\{\begin{array}[]{l}\nu_{1}=0.9\nu,\;\beta_{1}=0.5\\ \nu_{2}=0.1\nu,\;\beta_{2}=7.5\end{array}\right.\end{array}\right.$ (23)
with
$\nu=\nu_{1}+\nu_{2}=\mu^{{\rm NEq}}=0.5,1,2\;\rm{MPa}$ (24)
in the non-equilibrium stored-energy function (10)2. As labeled in (23), these material constants yield non-equilibrium stored-energy functions (10)2 with weaker ($\beta_{2}<\alpha_{2}$), equal ($\beta_{2}=\alpha_{2}$), and stronger ($\beta_{2}>\alpha_{2}$) growth conditions — in the limit of large deformations as $||{\bf F}||=\sqrt{I_{1}}\nearrow+\infty$ — than that of the equilibrium stored-energy function (10)1 with (22).
So as to also thoroughly probe the effects of the value of the viscosity and its nonlinearity, we consider the three sets of material constants
$\left\{\begin{array}[]{rl}\texttt{Constant Viscosity:}&\left\{\begin{array}[]{l}N_{0}=\eta_{0}\\ \eta_{\infty}=\eta_{0}\end{array}\right.\vspace{0.18cm}\\ \texttt{Shear-thinning Viscosity:}&\left\{\begin{array}[]{l}N_{0}=\eta_{0}\\ \eta_{\infty}=0.01\eta_{0}\\ K_{2}=500\,{\rm MPa}^{-2},\;\gamma_{2}=1\end{array}\right.\vspace{0.18cm}\\ \texttt{Deformation-dependent}&\vspace{-0.3cm}\\ \texttt{Viscosity:}&\left\{\begin{array}[]{l}N_{0}=100\eta_{0}\\ K_{1}=2,\;\gamma_{1}=4\\ K_{2}=0\end{array}\right.\end{array}\right.$ (25)
with
$\eta_{0}=\tau_{0}\nu=0.5,1,2\;\rm{MPa}\,{\rm s},\quad\tau_{0}=1\,{\rm s},$ (26)
in the viscosity function (13). As labeled in (25), these values serve to examine the basic case of constant viscosity, as well as those of shear-thinning and deformation-dependent viscosities, one at a time. Note that the value (26) of the initial viscosity $\eta_{0}$ is chosen in a normalized manner so that the resulting initial relaxation time is kept constant at $\tau_{0}=\eta_{0}/\mu^{{\rm NEq}}=1$ s for all the non-equilibrium initial shear moduli (24).
For the above material constants and pre-existing crack and specimen geometries, the entire range of viscoelastic behaviors — from elasticity-dominated to viscosity-dominated — occurs effectively within the range $\dot{\mathrm{\Lambda}}_{0}=[10^{-3},10^{3}]$ s-1 of global stretch rates. Accordingly, in the parametric study presented in this section, we carry out simulations for the nine global stretch rates
$\dot{\mathrm{\Lambda}}_{0}=0+,10^{-3},10^{-2},10^{-1},10^{0},10^{1},10^{2},10^{3},+\infty\;{\rm s}^{-1},$
which include the purely elastic limit of infinitesimally slow loading, $\dot{\mathrm{\Lambda}}_{0}=0+$, as well as the “pseudo-elastic” limit of infinitely fast loading, $\dot{\mathrm{\Lambda}}_{0}=+\infty$.
In all, there are $3\times 3\times 3=27$ combinations of material constants, $3$ crack lengths, and $9$ different stretch rates, which result in a total of $729$ simulations. In the following subsections, we present the main findings from these simulations, organized in terms of the types of viscosity of the elastomer and, within this, in terms of the growth conditions of its non-equilibrium elasticity and loading rate.
4.1 Constant viscosity
We begin by presenting results that illustrate how the critical values $\mathrm{\Lambda}_{c}$ and $S_{c}$ of the global stretch $\mathrm{\Lambda}$ and the global stress $S$, as well as the corresponding critical tearing energy $T_{c}$, at which the crack grows in single edge notch fracture tests depend on the loading rate and on the growth conditions of the non-equilibrium elasticity for the basic case when the viscosity of the elastomer is constant.
Figure 6 presents results for the global stress $S=P/(HB)$ as a function of the applied global stretch $\mathrm{\Lambda}=l/L$ for specimens stretched at the constant global stretch rates $\dot{\mathrm{\Lambda}}_{0}=10^{-3},10^{-1},10^{0}$, $10^{1},10^{3}$ s-1. The results pertain to specimens with initial crack size $A=2$ mm, elastomers with non-equilibrium elasticity of weaker, equal, and stronger growth conditions (23), with $\nu=\mu^{{\rm NEq}}=20\mu^{{\rm Eq}}=2$ MPa, than their equilibrium elasticity, and constant viscosity (25)1 with $\eta_{0}=\nu\tau_{0}=2$ MPa s. All the results are plotted up to the critical global stretch $\mathrm{\Lambda}_{c}$ at which the Griffith criticality condition (18) is satisfied (indicated by an empty circle in all the force-deformation plots) for the representative case when the critical energy release rate is $G_{c}=150$ N/m.
Figure 6: Force-deformation response of specimens with pre-existing cracks of length $A=2$ mm stretched at various constant global stretch rates $\dot{\mathrm{\Lambda}}_{0}$. The results pertain to elastomers with constant viscosity (25)1 and non-equilibrium elasticity of weaker (23)1, equal (23)2, and stronger (23)3 growth conditions than their equilibrium elasticity.
Figure 7: The critical global stretch $\mathrm{\Lambda}_{c}$, global stress $S_{c}$, and tearing energy $T_{c}$ at which the crack grows plotted as functions of the global stretch rate $\dot{\mathrm{\Lambda}}_{0}$ at which the test is carried out. The results pertain to elastomers with constant viscosity (25)1.
An immediate observation from the results in Fig. 6 is that the critical global stretches $\mathrm{\Lambda}_{c}$ and stresses $S_{c}$ at which the crack grows are dependent on the loading rate $\dot{\mathrm{\Lambda}}_{0}$ and on the growth conditions of the non-equilibrium elasticity of the elastomer, as characterized by the material constant $\beta_{2}$. Interestingly, the dependence of $\mathrm{\Lambda}_{c}$ on $\dot{\mathrm{\Lambda}}_{0}$ and $\beta_{2}$ is non-monotonic, while that of $S_{c}$ is monotonic on $\dot{\mathrm{\Lambda}}_{0}$ but not so on $\beta_{2}$.
The quantitative dependence of $\mathrm{\Lambda}_{c}$ and $S_{c}$ on $\dot{\mathrm{\Lambda}}_{0}$ and $\beta_{2}$ are better illustrated by Fig. 7. The results for the stretch show that, save for the special case of equal growth ($\beta_{2}=\alpha_{2}=2.5$), the smallest $\mathrm{\Lambda}_{c}$ is attained at the slowest loading rate $\dot{\mathrm{\Lambda}}_{0}=0+$, when the behavior of the material is purely elastic. As the loading rate $\dot{\mathrm{\Lambda}}_{0}$ increases so does $\mathrm{\Lambda}_{c}$, but only until certain threshold $\dot{\mathrm{\Lambda}}_{0}^{{\rm th}}$ at which $\mathrm{\Lambda}_{c}$ reaches a local maximum. Further increase in $\dot{\mathrm{\Lambda}}_{0}$ beyond that threshold leads to the evolution of $\mathrm{\Lambda}_{c}$ towards an asymptotic value. The threshold $\dot{\mathrm{\Lambda}}_{0}^{{\rm th}}$ at which $\mathrm{\Lambda}_{c}$ reaches its local maximum appears to correlate with the initial relaxation time $\tau_{0}=\eta_{0}/\mu^{{\rm NEq}}$, roughly, $\dot{\mathrm{\Lambda}}_{0}^{{\rm th}}\approx\tau_{0}^{-1}$. On the other hand, the results for the stress show that $S_{c}$ increases monotonically with increasing loading rate $\dot{\mathrm{\Lambda}}_{0}$, this irrespective of the value of $\beta_{2}$. The dependence of both $\mathrm{\Lambda}_{c}$ and $S_{c}$ on the non-equilibrium elasticity of the elastomer is clearly non-monotonic and intricately coupled with the loading rate.
In addition to the results for the critical global stretch $\mathrm{\Lambda}_{c}$ and the critical global stress $S_{c}$, it is instructive to plot the results for the corresponding critical tearing energy (21). Figure 7 includes such a plot in terms of the global stretch rate $\dot{\mathrm{\Lambda}}_{0}$ for the three material constants $\beta_{2}=1,2.5,7.5$. The results show the typical “S” shape found in other types of fracture tests, such as the trousers fracture testGreensmith and Thomas (1955), with the caveat that they are not strictly monotonic. Specifically, as $\dot{\mathrm{\Lambda}}_{0}\searrow 0$, for sufficiently slow global stretch rates, $T_{c}\searrow G_{c}$. As $\dot{\mathrm{\Lambda}}_{0}$ increases, so does $T_{c}$, although not necessarily monotonically. As $\dot{\mathrm{\Lambda}}_{0}\nearrow+\infty$, for sufficiently fast global stretch rates, $T_{c}$ approaches an asymptotic value, $T_{c}^{\infty}$ say. The transition of $T_{c}$ from its minimum value $G_{c}$ to its asymptotic value $T_{c}^{\infty}$ is controlled by both the non-equilibrium elasticity of the elastomer and its viscosity. Given that $G_{c}=150$ N/m and $T_{c}^{\infty}>3000$ N/m, the results in Fig. 7 make it plain that the critical tearing energy $T_{c}$ is primarily a direct manifestation of the viscoelastic behavior of the elastomer, and not of its fracture behavior. In other words, in view of the representation (3), the value of $T_{c}$ is mostly a measure of how much energy is dissipated by viscous deformation and not by the creation of surface.
4.2 Shear-thinning viscosity
Figure 8: The critical global stretch $\mathrm{\Lambda}_{c}$, global stress $S_{c}$, and tearing energy $T_{c}$ at which the crack grows plotted as functions of the global stretch rate $\dot{\mathrm{\Lambda}}_{0}$ at which the test is carried out. The results pertain to elastomers with shear-thinning viscosity (25)2.
Figure 9: The critical global stretch $\mathrm{\Lambda}_{c}$, global stress $S_{c}$, and tearing energy $T_{c}$ at which the crack grows plotted as functions of the global stretch rate $\dot{\mathrm{\Lambda}}_{0}$ at which the test is carried out. The results pertain to elastomers with deformation-dependent viscosity (25)3.
As already pointed out above, elastomers are by and large characterized not by a constant viscosity but by a strongly nonlinear viscosity that depends both on the deformation and on the deformation rate. In this subsection, we examine the effects of the latter on $\mathrm{\Lambda}_{c}$, $S_{c}$, and $T_{c}$.
Figure 8 presents results analogous to those presented in Fig. 7 for elastomers whose viscosity is not constant but instead decreases with deformation rate, that is, a viscosity that exhibits shear thinning; the corresponding force-deformation results analogous to those in Fig. 6 are presented in Fig. S2 of the Supplementary Material. Precisely, the results pertain to specimens with initial crack size $A=2$ mm, elastomers with non-equilibrium elasticity of weaker, equal, and stronger growth conditions (23), with $\nu=\mu^{{\rm NEq}}=20\mu^{{\rm Eq}}=2$ MPa, than their equilibrium elasticity, and shear-thinning viscosity (25)2 with $\eta_{0}=\nu\tau_{0}=2$ MPa s. Much like the results in Fig. 7, the results in Fig. 8 correspond to the representative case when the critical energy release rate is $G_{c}=150$ N/m.
A quick glance suffices to recognize that the results in Fig. 8 for shear-thinning viscosity are qualitatively similar to those in Fig. 7 for constant viscosity. Quantitatively, Fig. 8 shows that the presence of shear thinning shifts the threshold $\dot{\mathrm{\Lambda}}_{0}^{{\rm th}}$ at which $\mathrm{\Lambda}_{c}$ reaches a local maximum to larger values. Shear thinning leads as well to larger differences in the values that $\mathrm{\Lambda}_{c}$ can take, significantly more so than the differences that result from variations in the growth conditions of the non-equilibrium elasticity of elastomers with constant viscosity. For instance, note that, roughly, $\mathrm{\Lambda}_{c}=1.35$ at $\dot{\mathrm{\Lambda}}_{0}=10^{-3}$ s-1 while $\mathrm{\Lambda}_{c}=2.6$ at $\dot{\mathrm{\Lambda}}_{0}=10^{1}$ s-1 for all growth conditions. By contrast, the presence of shear thinning appears to have a more moderate quantitative effect on the critical global stress $S_{c}$ and critical tearing energy $T_{c}$.
4.3 Deformation-dependent viscosity
Next, we examine the effects that a nonlinear viscosity that depends on deformation can have on $\mathrm{\Lambda}_{c}$, $S_{c}$, and $T_{c}$.
Figure 9 presents results analogous to those presented in Figs. 7 and 8 for elastomers whose viscosity increases with increasing deformation; the corresponding force-deformation results analogous to those in Fig. 6 are presented in Fig. S3 of the Supplementary Material. In particular, the results pertain to specimens with initial crack size $A=2$ mm, elastomers with non-equilibrium elasticity of weaker, equal, and stronger growth conditions (23), with $\nu=\mu^{{\rm NEq}}=20\mu^{{\rm Eq}}=2$ MPa, than their equilibrium elasticity, and deformation-dependent viscosity (25)3 with $\eta_{0}=\nu\tau_{0}=2$ MPa s. Again, much like the results in the two preceding subsections, the results in Fig. 9 correspond to the representative case when the critical energy release rate is $G_{c}=150$ N/m.
Qualitatively, the results in Fig. 9 for a nonlinear viscosity that increases with deformation are similar to those in Fig. 7 for constant viscosity. Quantitatively, two observations are called for when comparing Fig. 9 with Fig. 7. First, the presence of a nonlinear viscosity that increases with deformation shifts the threshold $\dot{\mathrm{\Lambda}}_{0}^{{\rm th}}$ at which $\mathrm{\Lambda}_{c}$ reaches a local maximum to smaller values. Second, on its own, contrary to a shear-thinning viscosity, a nonlinear viscosity that increases with deformation does not result in a range of values for $\mathrm{\Lambda}_{c}$ that are substantially different from those that result from a constant viscosity. The same is true for $S_{c}$ and $T_{c}$.
4.4 The classical formula (4) of Rivlin and Thomas vs. 3D full-field results
The classical formula (4) of Rivlin and ThomasRivlin and Thomas (1953) — together with the approximation (5) for the numerical factor $K(\mathrm{\Lambda})$ that emerged from the works of GreensmithGreensmith (1963) and LindleyLindley (1972) — is commonly used well beyond its original scope (see Subsection 1.1 above) in the experimental literature so as to estimate the values of the critical tearing energy $T_{c}$ from single edge notch fracture tests. In the sequel, we show how such an overextended use can lead to significant errors in the estimation of $T_{c}$. We do so by confronting the formula (4) directly with an exact solution from our 3D full-field simulations for viscoelastic elastomers.
Remark 3
In order to use the formula (4) for elastomers that are not purely elastic solids but solids that dissipate energy by deformation, such as viscoelastic elastomers, the stored-energy function $W_{{\rm un}}(\mathrm{\Lambda})$ is typically reinterpreted as the “strain energy density” (stored and dissipated) of the elastomer when subjected to a uniaxial tension test that is carried out at the same stretch rate $\dot{\mathrm{\Lambda}}_{0}$ as the single edge notch fracture test of interest.
Figure 10 provides a comparison between the energy release rate $-\partial\mathcal{W}/\partial\mathrm{\Gamma}_{0}$ generated by the classical formula (4), with the numerical factor (5), and the exact result (21) determined from 3D full-field simulations of a single edge notch test, carried out at the global stretch rate $\dot{\mathrm{\Lambda}}_{0}=10^{0}$ s-1, for a specimen with pre-existing crack of length $A=2$ mm and an elastomer with shear-thinning viscosity (25)2, $\eta_{0}=\nu\tau_{0}=2$ MPa s, and non-equilibrium elasticity of equal growth conditions (23)2, with $\nu=\mu^{{\rm NEq}}=20\mu^{{\rm Eq}}=2$ MPa, than its equilibrium elasticity.
The comparison in Fig. 10 makes it clear that the classical formula (4) of Rivlin and ThomasRivlin and Thomas (1953) can be greatly inaccurate, increasingly more so for larger global stretches $\mathrm{\Lambda}$, when used for elastomers that dissipate energy by deformation and for cracks that are not much smaller than the length and the width of the specimen.
Figure 10: Representative comparison between the energy release rate generated by the classical Rivlin-Thomas formula (4), with the numerical factor (5), and the exact result (21) determined from 3D full-field simulations of a single edge notch test carried out at the global stretch rate $\dot{\mathrm{\Lambda}}_{0}=10^{0}$ s-1. The results pertain to a specimen with pre-existing crack of length $A=2$ mm and an elastomer with shear-thinning viscosity (25)2 and non-equilibrium elasticity of equal growth conditions (23)2 to its equilibrium elasticity.
5 Comparisons with experiments on an acrylate elastomer
As alluded to in the Introduction, since the pioneering work of BusseBusse (1934) in the 1930s, rather remarkably, only a handful of the numerous single edge notch fracture tests that have been reported in the literature have been carried out at different loading rates $\dot{\mathrm{\Lambda}}_{0}$. In this section, we deploy the Griffith criticality condition (2) to explain one of them, those of Le MennMenn (2022) on an acrylate elastomer.
As part of their doctoral thesis, Le MennMenn (2022) carried out a series of single edge notch tests on dog-bone specimens of gauge length $L=15$ mm, width $H=4$ mm, thickness $B\in[0.2,0.8]$ mm, and crack size $A=1$ mm made of acrylate elastomers (for use as coatings of optical fibers) at room temperature and four different constant global stretch rates, approximately given by $\dot{\mathrm{\Lambda}}_{0}=6.7\times 10^{-5},6.7\times 10^{-4},6.7\times 10^{-3},3.3\times 10^{-2}$ s-1; see Appendix B.4 in their work. Out of all the various compositions that Le MennMenn (2022) considered, we focus here on the one that she labeled OL HEA 30% PEA for which she provided experimental data for the elastic response of the elastomer in the form of stress-stretch data for uniaxial and equi-biaxial tension carried out at a constant stretch rate of about $\dot{\lambda}_{0}=6.7\times 10^{-3}$ s-1, sufficiently slow that viscous effects were not observed; see Chapter 3 in their work.
Figure 11: Comparison between the stress-stretch response described by the viscoelastic model (11)-(12), with the material constants in Table 1, and the experimental data reported by Le MennMenn (2022) for the acrylate elastomer OL HEA 30% PEA subjected to uniaxial and equi-biaxial tension applied at a slow constant stretch rate $\dot{\lambda}_{0}$ at room temperature.
Figure 12: (a) Comparison between the force-deformation response predicted theoretically and the experimental data reported by Le MennMenn (2022) for single edge notch fracture tests on the acrylate elastomer OL HEA 30% PEA carried out at four different global stretch rates $\dot{\mathrm{\Lambda}}_{0}$. (b, c) Corresponding comparison between the critical global stretch $\mathrm{\Lambda}_{c}$ and the critical global stress $S_{c}$ at which fracture nucleates predicted by the Griffith criticality condition (2) and the experimental results.
The viscoelastic behavior.
The fitting of the material constants $\mu_{1}$, $\mu_{2}$, $\alpha_{1}$, $\alpha_{2}$, which, again, describe the equilibrium elasticity in the viscoelastic model (11)-(12), to the uniaxial and equi-biaxial data reported by Le MennMenn (2022) yields the values listed in the first column of Table 1. As seen from the comparisons presented in Fig. 11, the constitutive relation (11)-(12) with such material constants describes reasonably well the experimentally measured response at states of thermodynamic equilibrium.
Table 1: Values of the material constants in the viscoelastic model (11)-(12) with (27) for the acrylate elastomer OL HEA 30% PEA.
[…]
$\alpha_{1}=4$ $\beta_{1}=-0.5$ $\eta_{\infty}=0$
[…]
$\alpha_{2}=1.9985$ $\beta_{2}=0$ $\gamma_{2}=1$
It remains to determine the materials constants $\nu_{1}$, $\nu_{2}$, $\beta_{1}$, $\beta_{2}$ for the non-equilibrium elasticity and the viscosity function $\eta({\bf C},{\bf C}^{v})$. Given the lack of direct experimental data beyond that presented in Fig. 11 for the viscoelastic behavior of OL HEA 30% PEA, we turn to remarking that the force-deformation response of the specimens in the single edge notch fracture tests reported by Le MennMenn (2022), reproduced in Fig. 12(a), differ only on the value of the critical global stretch $\mathrm{\Lambda}_{c}$ — and hence also on the value of the critical global stress $S_{c}$ — at which fracture nucleates. In other words, the force-deformation response of the specimens appears to be independent of the loading rate $\dot{\mathrm{\Lambda}}_{0}$, except for the fact that $\mathrm{\Lambda}_{c}$ and $S_{c}$ increase with increasing $\dot{\mathrm{\Lambda}}_{0}$. Given that the stretch and stretch rates around the crack front in single edge notch fracture tests are locally very large, such a behavior may be attributed to a viscosity that is strongly nonlinear at large stretches and stretch rates, but one that is negligible otherwise. With this mechanism in mind, we consider a viscosity function of the piecewise form
$\eta({\bf C},{\bf C}^{v})=\left\{\begin{array}[]{ll}\eta_{0},&\,\textbf{X}\in\mathrm{\Omega}_{0}^{\rm I}\vspace{0.2cm}\\ \eta_{\infty}+\dfrac{\eta_{0}-\eta_{\infty}}{1+\left(K_{2}\mathcal{J}_{2}^{\rm{NEq}}\right)^{\gamma_{2}}},&\,\textbf{X}\in\mathrm{\Omega}_{0}^{\rm II}\vspace{0.2cm}\\ \eta_{\infty},&\,\textbf{X}\in\mathrm{\Omega}_{0}\setminus(\mathrm{\Omega}_{0}^{\rm I}\cup\mathrm{\Omega}_{0}^{\rm II})\end{array}\right.$ (27)
where, for definiteness, $\mathrm{\Omega}_{0}^{\rm I}=\{{\bf X}:|X_{1}|\leq B/2,\,(X_{2}-X_{2}^{\rm I})^{2}+X_{3}^{2}<(0.05A)^{2}\}$ and $\mathrm{\Omega}_{0}^{\rm II}=(\{{\bf X}:|X_{1}|\leq B/2,\,(X_{2}-X_{2}^{\rm II})^{2}+X_{3}^{2}<(0.7A)^{2}\})\setminus\mathrm{\Omega}_{0}^{\rm I}$ with $X_{2}^{\rm I}=1.05A-0.05\,{\rm mm}$ and $X_{2}^{\rm II}=1.7A-0.05\,{\rm mm}$. As illustrated by Fig. 13, $\mathrm{\Omega}_{0}^{\rm I}$ stands for a small region surrounding the crack front where the viscosity is assumed to be of constant value $\eta_{0}$. In such a region, the stretch and stretch rates are extremely large and we assume that whatever shear-thinning of the viscosity may exist, it is canceled out by its enhancement due to deformation. Furthermore, $\mathrm{\Omega}_{0}^{\rm II}$ stands for a larger but still relatively small enclosing region where the viscosity is assumed to decrease from $\eta_{0}$ to $\eta_{\infty}\ll\eta_{0}$ due to shear thinning. Outside $\mathrm{\Omega}_{0}^{\rm I}$ and $\mathrm{\Omega}_{0}^{\rm II}$, in the remaining of the specimen, the viscosity is assumed to be of constant value $\eta_{\infty}$.
Given the viscosity function (27), a search of the material constants $\nu_{1}$, $\nu_{2}$, $\beta_{1}$, $\beta_{2}$, $\eta_{0}$, $\eta_{\infty}$, $K_{2}$, $\gamma_{2}$ in the viscoelastic model (11)-(12) that are consistent with the data of Le MennMenn (2022) yields the values listed in the second and third columns of Table 1.
Figure 13: Schematics of the small regions $\mathrm{\Omega}_{0}^{\rm{I}}$ and $\mathrm{\Omega}_{0}^{\rm{II}}$ around the crack front in the single edge notch fracture tests carried out by Le MennMenn (2022) wherein the viscoelastic behavior of the acrylate elastomer OL HEA 30% PEA is described by the nonlinear viscosity (27)1,2.
The critical energy release rate.
In addition to the viscoelastic behavior of the elastomer, the use of the Griffith criticality condition (2) requires knowledge of its critical energy release rate $G_{c}$. For the problem at hand, the fitting of $G_{c}$ in (2) from the single edge notch fracture test carried out at the slowest loading rate $\dot{\mathrm{\Lambda}}_{0}=6.7\times 10^{-5}$ s-1 yields
$G_{c}=92\;{\rm N}/{\rm m},$ (28)
which is consistent with the estimate provided by Le MennMenn (2022); see Table 3.7 in their work.
At this stage, having calibrated the viscoelastic behavior and the critical energy release rate of the elastomer, we are ready to deploy the Griffith criticality condition (2) to confront its predictions with the single edge notch fracture tests reported by Le MennMenn (2022). Figures 12(b,c) do so for the critical global stretch $\mathrm{\Lambda}_{c}$ and the critical global stress $S_{c}$ at which fracture nucleates as functions of the applied global stretch rate $\dot{\mathrm{\Lambda}}_{0}$. The error bars for the experimental data correspond to the neighborhood over which the force-deformation response of the specimens approaches its peak value and hence over which the onset of fracture is bound to occur.
The main observation from Figs. 12(b, c) is that the Griffith criticality condition (2) is in good overall agreement with the experimental observations. There is an appreciable quantitative difference between the theoretically predicted values for the critical global stress $S_{c}$ and those measured experimentally. This difference is consistent with that shown by Fig. 12(a) between the force-deformation response of the specimens predicted by the theory and that measured experimentally. To wit, the response in the experiments is softer from the outset. This softer behavior may be due to the cracks in the experiments being actually larger than $A=1$ mm and/or to some experimental error in the measurement of the global stretch $\mathrm{\Lambda}$. Howbeit, both sets of results exhibit essentially the same behavior in the force-deformation response and in the critical global stretch $\mathrm{\Lambda}_{c}$ and critical global stress $S_{c}$ at fracture.
6 Summary and final comments
The parametric study presented in Section 4 and the comparisons with experiments presented in Section 5 have shown that the nonlinear viscoelastic properties of elastomers have a profound effect on their behavior in single edge notch fracture tests. Precisely, the critical global stretch $\mathrm{\Lambda}_{c}$ and the critical global stress $S_{c}$ at which the pre-existing crack in such tests starts to grow depend on the non-Gaussian elasticity of the elastomer at hand and, even more so, on the nonlinearity of its viscosity. For tests carried out at constant global stretch rates $\dot{\mathrm{\Lambda}}_{0}$, the smallest values of the critical global stretch $\mathrm{\Lambda}_{c}$ and the critical global stress $S_{c}$ are typically attained at infinitesimally slow loading rates $\dot{\mathrm{\Lambda}}_{0}=0+$, when viscous effects are absent. As the loading rate increases, so does $\mathrm{\Lambda}_{c}$, but not necessarily in a monotonic fashion. On the other hand, $S_{c}$ appears to always increase monotonically with increasing loading rate. Elastomers with shear-thinning viscosity tend to exhibit the strongest sensitivity to the loading rate.
The strong sensitivity of the results of single edge notch fracture tests on the nonlinear viscoelastic properties of elastomers is due to the fact that the region around the crack front in such tests experiences a wide range of very large stretches and stretch rates — more so than in other standard fracture tests, such as the “pure-shear” and the trousers tests — that fully probes the non-Gaussian elasticity and the nonlinear viscosity of the elastomer being tested. As a corollary of this sensitivity, one can envision using single edge notch fracture tests not to measure fracture properties but to indirectly measure viscoelastic properties at large stretches and stretch rates that would be otherwise inaccessible by direct experimental techniques.
Additionally, the parametric study presented in Section 4 has served to establish that the classical formula (4) of Rivlin and ThomasRivlin and Thomas (1953), together with the approximation (5) for the numerical factor $K(\mathrm{\Lambda})$ due to GreensmithGreensmith (1963) and LindleyLindley (1972), can be greatly inaccurate when used to estimate the critical tearing energy $T_{c}$ for elastomers that dissipate energy by deformation and for cracks that are not much smaller than the length and the width of the specimen. This result should not come as a surprise, since Rivlin and ThomasRivlin and Thomas (1953) derived the formula (4) precisely under the premise that energy dissipation (viscous or otherwise) by deformation is absent and that the pre-existing crack is much smaller than the length and the width of the specimen, but much larger than its thickness.
The comparisons with experiments presented in Section 5 have also served to add to previous validation results for other tests and materialsShrimali and Lopez-Pamies (2023a, b, c) and thus they have provided further evidence that the Griffith criticality condition (2) may indeed be the condition that governs when large cracks grow in viscoelastic elastomers subjected to arbitrary quasi-static loading conditions.
To conclusively establish whether the Griffith criticality condition (2) is “the law of the land”, more comparisons with experiments on different elastomers under different types of loadings would be essential. In this regard, we emphasize that a key implication of (2) is that any rate effects on the nucleation of fracture from large pre-existing cracks in elastomers are due to their nonlinear viscoelastic behavior. This, we hope, will motivate experimentalists and theoreticians alike to thoroughly characterize the nonlinear viscoelasticity of elastomers over large ranges of stretches and stretch rates — and not just within the classical limited setting of small deformations, as most often donePersson et al. (2024), which is insufficient to analyze fracture — so as to help establishing whether all rate effects observed in the fracture of elastomers are indeed of a viscous nature.
Acknowledgements
This work was supported by the National Science Foundation (USA) through the Grant DMS–2308169. This support is gratefully acknowledged.